# Joint parameter estimations for spin glasses

### v1

@@ 1. Introduction and Main Results
$P_{\beta,h}({\sigma})=\frac{1}{Z_{N}(\beta,h)}\exp(\beta\langle{\sigma},J{\sigma}\rangle+h\langle{\sigma},\boldsymbol{1}\rangle),$
@@
In this paper, we aim to show that MPLE is a $\sqrt{N}$-jointly consistent estimator for $(\beta,h)$ allowing $J$ to possess positive and negative entries. More precisely, the matrix $J$ is a weighted adjacency matrix of a graph, where the weights are i.i.d. centered random variable with finite third moment and the graph requires that its averaging degree diverges at least logarithmically fast and is comparable to the maximum degree. This setting covers a collection of spin glass models such as the celebrated SK model and its diluted variants. We show that the validity of our result requires checking three major conditions: the non-flatness of the free energy, the boundedness of the operator norm of $J$, and the positivity of the Hessian matrix of the pseudolikelihood function, see Theorem 1.1. Among them, the third condition plays a central role and it also appeared in [GM20]. Incidentally, the authors therein verified this positivity condition critically relying on the nonnegativity assumption of coupling coefficients which allowed them to use the nonlinear large deviation theory for random graphs [CD16, BM17, EG18, Chat17]. However, such technique does not seem to work when the coupling matrix $J$ has both positive and negative entries. Until now, it remained as a main obstacle to extend joint estimation results beyond the non-negative coupling coefficients setting. We are able to check the non-singularity condition with high probability in the setting of random coupling matrices via the small ball probability arguments [RV08, Ver14].
@@ 1.1. Notation
Let $\mathcal{G}:=([N],E)$ be a graph with vertex set $[N]=\{1,2,\ldots,N\}$ and edge set $E$. Denote by $d_{i}$ and $d$ the degree of vertex $i\in[N]$ and the average degree of $\mathcal{G}$, respectively. In addition, $d_{\max}:=\max_{i\in[N]}d_{i}$ denotes the maximum degree of $\mathcal{G}$. Let $G=(g_{ij})_{i,j=1}^{N}$ be a symmetric matrix and vanishes along the diagonal, i.e., $g_{ii}=0$ for $i\in[N]$. Set
@@ 1.2. Basic Setting of MPLE
$H(\beta,h|{\sigma}):=\begin{pmatrix}&\sum_{i=1}^{N}m_{i}({\sigma})^{2}\theta_{i}(\beta,h|{\sigma})^{2}&\sum_{i=1}^{N}m_{i}({\sigma})\theta_{i}(\beta,h|{\sigma})^{2}\\ &\sum_{i=1}^{N}m_{i}({\sigma})\theta_{i}(\beta,h|{\sigma})^{2}&\sum_{i=1}^{N}\theta_{i}(\beta,h|{\sigma})^{2}\end{pmatrix},$ (7)
@@ 1. $(i)$
    Non-flatness of the limiting free energy :::: There exists some $0<\beta^{\prime}<\beta_{0}$ such that
    $\liminf_{N\to\infty}(F_{N}(\beta^{\prime},h_{0})-F_{N}(0,h_{0}))>0.$
@@ 2. $(ii)$
    Boundedness of $J$::::
    $\|J\|=O(1),$
    where $\|J\|$ is the operator norm of $J$.
  3. $(iii)$
    Positivity of the Hessian :::: Under $P_{\beta_{0},h_{0}}$,
@@ 3. $(iii)$
Note that this theorem holds for arbitrary real-valued symmetric matrices $J$ with zeros along the diagonal and they do not need to be of the form (1).
@@
Under different assumptions, Ghosal-Mukherjee [GM20] obtained similar existence and consistency results as Theorem 1.2, where the matrix $J$ has nonnegative entries and uniformly bounded row sums. Our proof for the consistency part in Theorem 1.1 is similar to that of [GM20]*Theorem 1.2. However, the approach for the existence part in the same theorem does not apply to our setting.
@@ 2. $(ii)$
Heuristically, the above condition specifies that the underlying graph can neither be too sparse nor have some vertices with very large degrees. Recall that a random variable $X$ is $\kappa^{2}$-subgaussian if $\mathbb{E}X=0$ and $\mathbb{E}e^{sX}\leqslant e^{\kappa^{2}s^{2}/2}$ for all $s\in\mathbb{R}.$ We say that $G=(g_{ij})_{i,j\in[N]}$ is a symmetric $\kappa^{2}$-subgaussian matrix with vanishing diagonal if its upper triangular entries are independent $\kappa^{2}$-subgaussian. Now we state the first result that verifies the non-flatness of the free energy in temperature.
Theorem 1.4.
Suppose that $G$ is a symmetric $\kappa^{2}$-subgaussian matrix with vanishing diagonal and Assumption 1.3 $(i)$ is in force. Then for any $\beta>0$ and $h\in\mathbb{R}$, there exist positive constants $c_{0},c,K$ independent of $N$ such that
$\mathbb{P}\bigl{(}F_{N}(\beta,h)-F_{N}(0,h)\leqslant c_{0}\bigr{)}\leqslant K\exp(-cN).$
The next result establishes the bounednedness of the operator norm $J$.
@@
As we shall see below, the proofs for these two theorems are not too difficult, which borrow some standard ideas and results from the treatment of the Ising model as well as the concentration inequalities for sparse random matrices [BvH16, CDG23]. The much more difficult component of this work lies on the verification of the strict positivity of the Hessian stated below.
@@
[…]
@@
As of now, with high probability, all three conditions in Theorem 1.1 have been verified under varying assumptions. Putting them all together, we obtain the asserted joint estimation for $(\beta,h).$
@@ 1.4. Proof Sketches
Theorem 1.1 consists of existence and consistency parts. The latter follows from a standard Delta method argument similar to [GM20]*Theorem 1.2. For existence, note that if the pseudolikelihood function diverges to the negative infinity as $|\beta|+|h|\to\infty$, then it has a unique maximizer (i.e., the MPLE exists) as it is already strictly concave thanks to the condition (iii). Almost the entirety of the proof of existence part is devoted to showing this divergence. We prove that this is indeed true if $(\tau_{i},m_{i}(\tau))_{i\in[N]}$ contains all possible signs. We then show that this property holds w.h.p. under the assumptions $(i)$ and $(ii)$. Even though this is intuitively clear since the spins and the fields $m_{i}$ should not be in favor of certain choices of signs overwhelmingly, proving it rigorously takes some effort. The key role is played by a general concentration result, see Lemma 2.1 below, which says that for $\tau\sim P_{\beta,h}$ and any suitably Lipschitz function $f$, the linear statistics $N^{-1}\sum_{i=1}^{N}f(\tau_{i},\tau_{i}m_{i}(\tau))$ stays around its conditional mean $N^{-1}\sum_{i=1}^{N}\langle f(\tau_{i},\tau_{i}m_{i}(\tau))|\tau_{l},\forall l\neq i\rangle.$ As an immediate consequence, our lemma implies the concentration for the score functions $S(\beta,h|\tau)$ and $Q(\beta,h|\tau)$ (see Corollary 2.2 below), which was previously established in [Chat07] using Stein’s method for exchangeable pairs. Our result holds for more general functions and its proof relies on the cavity argument instead.
For Theorem 1.4, it suffices to show that $\operatorname{\mathds{E}}F_{N}(\beta,h)$ is strictly increasing in $\beta>0$ since this quantity is concentrated. To this end, we first show that the expected free energy is universal in the sense that $\operatorname{\mathds{E}}F_{N}(\beta,h)$ is asymptotically the same between Gaussian and subgaussian matrices with proper matching moments. Now for the free energy with Gaussian disorder, one can use the Gaussian integration by parts and then a standard cavity computation to write
$\partial_{\beta}\operatorname{\mathds{E}}F(\beta,h)$ $=\frac{\beta^{2}}{2}\Bigl{(}1-\frac{2}{Nd}\sum_{(i,j)\in E}\operatorname{\mathds{E}}\langle\sigma_{i}^{1}\sigma_{j}^{1}\sigma_{i}^{2}\sigma_{j}^{2}\rangle\Bigr{)}$
$=\frac{\beta^{2}}{2}\Bigl{(}1-\frac{2}{Nd}\sum_{(i,j)\in E}\operatorname{\mathds{E}}\bigl{\langle}\tanh(\beta m_{i}(\sigma)+h)\sigma_{j}\bigr{\rangle}^{2}\Bigr{)}.$
The crucial observation that we will establish below is that with high probability, if ${\sigma}\sim P_{\beta,h}(\cdot)$, then there exists a vertex subset $U=U({\sigma})\subseteq[N]$ that is of linear size and the fields $(m_{i}({\sigma}))_{i\in U}$ are uniformly bounded by an absolute constant. Additionally, each $i\in U$ satisfies $d_{i}\geqslant d/4$. Using these, we readily see that $\partial_{\beta}\mathbb{E}F(\beta,h)\geqslant c$ for some universal constant $c$ as long as $\beta$ is small enough and yields non-flatness of $F_{N}(\beta,h)$. The proof of Theorem 1.5 follows straightforwardly by using the asymptotic results [BvH16] and the concentration bound [CDG23] for the operator norm of sparse subgaussian matrices.
The proof of Theorem 1.6 is the central ingredient of this work. For simplicity, let us sketch an argument to show that
@@
It is easy to check using Markov inequality that $\|{m}({\sigma})\|_{\infty}\leqslant M:=2^{N}$ with probability at least $1-C(1.5)^{-N}$. On this event, the infimum over $\gamma$ in (12) can be restricted to the interval $[-M,M]$, which can further be approximated by taking minima over an $\epsilon$-net $\mathcal{D}$ of size $O(2^{N}\varepsilon^{-1})$ for small but constant $\epsilon>0$. Therefore, it suffices to show that uniformly for any ${\sigma}\in\{-1,1\}^{N}$ and any $\gamma\in\mathcal{D}$,
@@
To overcome this obstacle, we borrow an idea from [Ver14]. We only use the randomness of a principal submatrix of $J$ of linear size while conditioning out the rest of the entries. Let $T\subset[N]$ be the set of the row (and the column) indices belonging to this principal submatrix, which we denote by $J_{T}$. Dealing with this smaller matrix instead of $J$ has the critical advantage that the entries of $J_{T}$ are independent (as long as $|T|\leqslant N/2$). We can then apply the previous small ball probability argument for $J_{T}\sigma_{T}$. However, successfully implementing this strategy requires that the matrix $J_{T}$ is not too sparse in the sense that in the graph $\mathcal{G}$ the number of edges between each vertex of $T$ and $[N]\setminus T$ tends to $\infty$ as $N\to\infty$. We apply Erdös’ probabilistic method to guarantee the existence of such a ‘good’ deterministic set $T$ provided $\mathcal{G}$ satisfies Assumption 1.3$(i)$ and $d\to\infty$ as $N\to\infty$. This gives the lower bound for $\min_{{\sigma}}T_{N}({\sigma})$.
@@
In the order their statements appear, the proofs for Theorems 1.1, 1.4, 1.5, and 1.6 are accordingly located in Sections 2, 3, 4, and 5.
@@
The proof of Theorem 1.1 contains the existence and consistency of the MPLE. The latter is deferred to Section 2.3. For the existence part, as mentioned in the proof sketches, it suffices to show that the pseudolikelihood function diverges to negative infinity at infinity. We verify this property in Section 2.2 based on a number of lemmas established in the following subsection.
@@ 2.1. Auxiliary Lemmas
Let $\beta>0$ and $h\in\mathbb{R}$. Let $f(x,y)$ be defined on $\{-1,1\}\times\mathbb{R}$. For any $i\in[N]$ and ${\sigma}\in\Sigma_{N},$ set
$L_{i}({\sigma})$ $=\langle f(\tau_{i},\tau_{i}m_{i}(\tau))|\tau_{l}=\sigma_{l},\forall l\neq i\rangle$
$=\frac{e^{\beta m_{i}({\sigma})+h}f(1,m_{i}({\sigma}))+e^{-\beta m_{i}({\sigma})-h}f(-1,-m_{i}({\sigma}))}{e^{\beta m_{i}({\sigma})+h}+e^{-\beta m_{i}({\sigma})-h}},$
[…]
@@
Let ${\tau}$ be sampled from $P_{\beta,h}$ for some $\beta>0$ and $h\in\mathbb{R}$. Assume that $f(x,y)$ be defined on $\{-1,1\}\times\mathbb{R}$ and satisfies $\|f\|_{\infty}\leqslant 1,$ $\|\partial_{y}f\|_{\infty}<\infty$, and $\|\partial_{y}^{2}f\|_{\infty}<\infty.$ For any $N\geqslant 1,$ we have that
$\Bigl{\langle}\Bigl{(}\sum_{i=1}^{N}\bigl{(}f(\tau_{i},\tau_{i}m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\Bigr{)}^{2}\Bigr{\rangle}$ $\leqslant 4N+C_{1}N\|J\|+\frac{C_{2}}{2}N\|J\|^{2},$ (14)

### v2

@@ 1. Introduction and Main Results
$P_{\beta,h}({\sigma})=\frac{1}{Z_{N}(\beta,h)}\exp(\beta\langle{\sigma},J{\sigma}\rangle/2+h\langle{\sigma},\boldsymbol{1}\rangle),$
@@
In this paper, we aim to show that MPLE is a $\sqrt{N}$-jointly consistent estimator for $(\beta,h)$ allowing $J$ to possess positive and negative entries. More precisely, the matrix $J$ is a weighted adjacency matrix of a graph, where the weights are i.i.d. centered random variable with finite third moment and the graph requires that its averaging degree diverges at least logarithmically fast and is comparable to the maximum degree. This setting covers a collection of spin glass models such as the celebrated SK model and its diluted variants. We show that the validity of our result requires checking two major conditions: the boundedness of the operator norm of $J$, and the positivity of the Hessian matrix of the pseudolikelihood function, see Theorem 1.1. Among them, the second condition plays a central role and it also appeared in [GM20]. Incidentally, the authors therein verified this positivity condition critically relying on the nonnegativity assumption of coupling coefficients which allowed them to use the nonlinear large deviation theory for random graphs [CD16, BM17, EG18, Chat17]. However, such technique does not seem to work when the coupling matrix $J$ has both positive and negative entries. Until now, it remained as a main obstacle to extend joint estimation results beyond the non-negative coupling coefficients setting. We are able to check the non-singularity condition with high probability in the setting of random coupling matrices via the small ball probability arguments [RV08, Ver14].
@@ 1.1. Notation
Let $\mathcal{G}:=([N],E)$ be a graph with vertex set $N]=\{1,2,\ldots,N\}$ and edge set $E$. Denote by $d_{i}$ and $d$ the degree of vertex $i\in[N]$ and the average degree of $\mathcal{G}$, respectively. In addition, $d_{\max}:=\max_{i\in[N]}d_{i}$ denotes the maximum degree of $\mathcal{G}$. Let $G=(g_{ij})_{i,j=1}^{N}$ be a symmetric matrix and vanishes along the diagonal, i.e., $g_{ii}=0$ for $i\in[N]$. Set
@@ 1.2. Basic Setting of MPLE
$H(\beta,h|{\sigma}):=\begin{pmatrix}&\sum_{i=1}^{N}m_{i}({\sigma})^{2}\theta_{i}(\beta,h|{\sigma})&\sum_{i=1}^{N}m_{i}({\sigma})\theta_{i}(\beta,h|{\sigma})\\ &\sum_{i=1}^{N}m_{i}({\sigma})\theta_{i}(\beta,h|{\sigma})&\sum_{i=1}^{N}\theta_{i}(\beta,h|{\sigma})\end{pmatrix}$, (7)
@@ 1. $(i)$
    Boundedness of $J$::
[…]
@@ 2. $(ii)$
    Positivity of the Hessian :: Under $P_{\beta_{0},h_{0}}$,
@@
Note that this theorem holds for arbitrary real-valued symmetric matrices $J$ with zeros along the diagonal and they do not need to be of the form (1). We further remark that the results can be easily extended to include the regime $\beta_{0}<0$. For notational and technical convenience, we focus on the regime $\beta_{0}>0$ as stated above.
@@
Under different assumptions, Ghosal-Mukherjee [GM20] obtained similar existence and consistency results as Theorem 1.2, where the matrix $J$ has nonnegative entries and uniformly bounded row sums. Our proof for the consistency part in Theorem 1.1 is similar to that of Theorem 1.2 in [GM20]. However, the approach for the existence part in the same theorem does not apply to our setting.
@@
Heuristically, the above condition specifies that the underlying graph can neither be too sparse nor have some vertices with very large degrees. Ghosal and Mukherjee’s work focused on non-disordered systems, where their coupling matrix $J$ is defined as $J=\frac{1}{d}\cdot A$, with $A$ being the adjacency matrix of the underlying graph and $d$ the average degree. This contrasts with our setup, which employs a scaling factor of $\frac{1}{\sqrt{d}}$ and assigns independent, mean-zero random weights to each edge. Their paper operates under the standing assumption of row sum boundedness for $J$, implying $d_{max}\leqslant Cd,$ a condition we also assume. Furthermore, their results show that
  • (i)
    If $d=O(1)$, the MPLE is $\sqrt{N}$-consistent.
  • (ii)
    If $d\to\infty$, depending on whether the underlying graph is irregular or regular, the MPLE is either $\sqrt{N}$-consistent and the estimation is not possible respectively.
To compare with our results, we do not have any results in the regime $d=O(1)$ due to the our standing assumption $d\gg\log N$. Second, under assumption $d\gg\log N$, we do not need any conditions on the regularity of the underlying graph.
Recall that a random variable $X$ is $\kappa^{2}$-subgaussian if $\mathbb{E}X=0$ and $\mathbb{E}e^{sX}\leqslant e^{\kappa^{2}s^{2}/2}$ for all $s\in\mathbb{R}.$ We say that $G=(g_{ij})_{i,j\in[N]}$ is a symmetric $\kappa^{2}$-subgaussian matrix with vanishing diagonal if its upper triangular entries are independent $\kappa^{2}$-subgaussian. Now we state the first result that verifies the boundedness of the operator norm of $J$.
@@
As we shall see below, the proofs for the above theorem are not too difficult, which borrow some standard ideas and results from the concentration inequalities for sparse random matrices [BvH16, CDG23]. The much more difficult component of this work lies on the verification of the strict positivity of the Hessian stated below.
@@
$\text{$(g_{ij})_{i>j}$ are independent with}\ \operatorname{\mathds{E}}{g_{ij}}=0,\,{\operatorname{\mathds{E}}[g_{ij}^{2}]}=1,\,\text{and}\,\operatorname{\mathds{E}}|g_{ij}|^{3}\leqslant B,$ (11)
@@
As of now, with high probability, all conditions in Theorem 1.1 have been verified under varying assumptions. Putting them all together, we obtain the asserted joint estimation for $(\beta,h).$
@@ 2. $(2)$
Remark 1.9.
It was indicated in [GM20] that the MPLE estimators are not jointly consistent when $J$ is the scaled adjacency matrix of any asymptotically regular graph $\mathcal{G}$ with the average degree diverging to infinity. In particular, consider the case when $\mathcal{G}=G(N,p_{N})$ is a dense Erdös-Rényi random graph with $p_{N}=p$ for some fixed $p\in(0,1]$. For $t>0$, let $\Theta_{t}=\{(\beta,h)\in(0,\infty)^{2}:t=\tanh(\beta t+h)\}.$ It was shown in [GM20, Theorem 1.6] that for any $(\beta_{0},h_{0})\in\Theta_{t}$, the measure $\mathbb{Q}_{n}\times\mathcal{G}$ is contiguous to $P_{\beta_{0},h_{0}}\times\mathcal{G}$, where $\mathbb{Q}_{n}$ is the product of Rademacher measures on $\{-1,1\}^{N}$ with mean $\tanh(t)$. Now, note that for any $(\beta_{0},h_{0})\in(0,\infty)^{2}$, there exists a unique $t\in(0,1)$ that satisfies $t=\tanh(\beta_{0}t+h_{0})$ and that one can always find a distinct $(\beta_{1},h_{1})\in\Theta_{t}.$ This statement readily implies that there does not exist any estimators that are jointly consistent for any $(\beta_{0},h_{0})\in(0,\infty)^{2}$. However, for the same interaction graph $\mathcal{G}$, if $J$ is of the form as (1), then we know from Theorem 1.7 that the MPLE are jointly consistent for any $\beta_{0}>0$ and $h_{0}\in\mathbb{R}$. This example suggests that the estimation problem when $J$ has only non-negative entries is significantly different from the case when $J$ is allowed to have both positive and negative entries.
@@ 1.4. Proof Sketches
Theorem 1.1 consists of existence and consistency parts. The latter follows from a standard Delta method argument similar to [GM20, Theorem 1.2]. For existence, note that if the pseudolikelihood function diverges to the negative infinity as $|\beta|+|h|\to\infty$, then it has a unique maximizer (i.e., the MPLE exists) as it is already strictly concave thanks to the condition $(ii)$ in Theorem 1.1. Almost the entirety of the proof of existence part is devoted to showing this divergence. We prove that this is indeed true if $(\tau_{i},m_{i}(\tau))_{i\in[N]}$ contains enough different signs. More precisely, the sufficient condition we need is stated in Lemma 2.3 and we then show that this property holds w.h.p. under the assumptions $(i)$ and $(ii)$. Even though this is intuitively clear since the spins and the fields $m_{i}$ should not be in favor of certain choices of signs overwhelmingly, proving it rigorously takes some effort. The key role is played by a general concentration result, see Lemma 2.1 below, which says that for $\tau\sim P_{\beta,h}$ and any suitably Lipschitz function $f$, the linear statistics $N^{-1}\sum_{i=1}^{N}f(\tau_{i},m_{i}(\tau))$ stays around its conditional mean $N^{-1}\sum_{i=1}^{N}\langle f(\tau_{i},m_{i}(\tau))|\tau_{l},\forall l\neq i\rangle.$ As an immediate consequence, our lemma implies the concentration for the score functions $S(\beta,h|\tau)$ and $Q(\beta,h|\tau)$ (see Corollary 2.2 below), which was previously established in [Chat07] using Stein’s method for exchangeable pairs. Our result holds for more general functions and its proof relies on the cavity argument instead.
The proof of Theorem 1.4 follows straightforwardly by using the asymptotic results [BvH16] and the concentration bound [CDG23] for the operator norm of sparse subgaussian matrices. The proof of Theorem 1.5 is the central ingredient of this work. For simplicity, let us sketch an argument to show that
@@
It is easy to check using Markov inequality that $\|{m}({\sigma})\|_{\infty}\leqslant M:=2^{N}$ with probability at least $1-C(1.5)^{-N}$. On this event, the infimum over $\gamma$ in (12) can be restricted to the interval [M,M][-M,M], which can further be approximated by taking minima over an $\epsilon$-net $\mathcal{D}$ of size $O(2^{N}\varepsilon^{-1})$ for small but constant $\epsilon>0$. Therefore, it suffices to show that uniformly for any ${\sigma}\in\{-1,1\}^{N}$ and any $\gamma\in\mathcal{D}$,
@@
To overcome this obstacle, we borrow an idea from [Ver14]. We only use the randomness of a principal submatrix of $J$ of linear size while conditioning out the rest of the entries. Let $T\subset[N]$ be the set of the row (and the column) indices belonging to this principal submatrix, which we denote by $J_{T}$. Dealing with this smaller matrix instead of $J$ has the critical advantage that the entries of $J_{T}$ are independent (as long as $|T|\leqslant N/2$). We can then apply the previous small ball probability argument for $J_{T}\sigma_{T}$. However, successfully implementing this strategy requires that the matrix $J_{T}$ is not too sparse in the sense that in the graph $\mathcal{G}$ the number of edges between each vertex of $T$ and $N]\setminus T$ tends to $\infty$ as $N\to\infty$. We apply Erdös’ probabilistic method to guarantee the existence of such a ‘good’ deterministic set $T$ provided $\mathcal{G}$ satisfies Assumption 1.3(i)(i) and $d\to\infty$ as $N\to\infty$. This gives the lower bound for $\min_{{\sigma}}T_{N}({\sigma})$.
@@
In the order their statements appear, the proofs for Theorems 1.1, 1.4, and 1.5 are accordingly located in Sections 2, 3, and 4.
@@
The proof of Theorem 1.1 contains the existence and consistency of the MPLE. The latter is deferred to Section 2.3. For the existence part, as mentioned in the proof sketches, it suffices to show that the pseudo-likelihood function diverges to negative infinity at infinity. We verify this property in Section 2.2 based on a number of lemmas established in the following subsection. We include a flowchart in Figure 1 for the high-level proof strategy of Theorem 1.1.
Lemma 2.5Lemma 2.1Lemma 2.4Lemma 2.3Corollary 2.2Theorem 1.1: ExistenceTheorem 1.1: Consistencyverifying conditions
Figure 1. Roadmap for the proof of Theorem 1.1. The dashed arrows represent lemma(s) used to validate hypothesis in the target lemma. The solid arrows present proof dependence.
@@ 2.1. Auxiliary Lemmas
Let $\beta>0$ and $h\in\mathbb{R}$. For $i\in[N],$ let $f_{i}(x,y)$ be defined on $\{-1,1\}\times\mathbb{R}$ and set
$L_{i}({\sigma})$ $=\langle f_{i}(\tau_{i},m_{i}(\tau))|\tau_{l}=\sigma_{l},\forall l\neq i\rangle$
$=\frac{e^{\beta m_{i}({\sigma})+h}f_{i}(1,m_{i}({\sigma}))+e^{-\beta m_{i}({\sigma})-h}f_{i}(-1,m_{i}({\sigma}))}{e^{\beta m_{i}({\sigma})+h}+e^{-\beta m_{i}({\sigma})-h}}$
for ${\sigma}\in\Sigma_{N},$ where $\langle\cdot\rangle$ is the Gibbs expectation with respect to $\tau\sim P_{\beta,h}.$ Our first lemma establishes the following concentration for the linear statistics of $(\tau_{i},m_{i}(\tau))_{i\in[N]}$.
@@
Let $\tau$} be sampled from $P_{\beta,h}$ for some $\beta>0$ and $h\in\mathbb{R}$. Let $\{f_{i}\}_{i\in[N]}$ be a collection of functions defined on $\{-1,+1\}\times\mathbb{R}$ with $\max_{i\in[N]}|f_{i}(x,y)|\leqslant C(1+|y|)$ for all $(x,y)\in\{-1,+1\}\times\mathbb{R}$ for some positive constant $C$ and $\max_{i\in[N]}(\|\partial_{y}f_{i}\|_{\infty},\|\partial_{y}^{2}f_{i}\|_{\infty})<\infty.$ Then there exists a constant $K$ depending only on $\beta$ and $C$ such that for any $N\geqslant 1,$ we have that
$\Bigl{\langle}\Bigl{(}\sum_{i=1}^{N}\bigl{(}f_{i}(\tau_{i},m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\Bigr{)}^{2}\Bigr{\rangle}$
$\leqslant K(1+\|J\|+\|J\|^{2})\Bigl{(}N+N\|J\|^{2}+\sum_{i=1}^{N}\|\partial_{y}f_{i}\|_{\infty}^{2}+\sum_{i=1}^{N}\|\partial_{y}^{2}f_{i}\|_{\infty}^{2}\Bigr{)}.$
Proof.
For any $i,j\in[N]$, set
$m_{j}^{(i)}({\sigma}):=\sum_{k\neq i,j}J_{jk}\sigma_{k}$
and let $L^{(i)}_{j}({\sigma})$ be the same as $L_{j}({\sigma})$ with $m_{j}({\sigma})$ replaced by $m_{j}^{(i)}({\sigma})$. Since $L_{i}(\tau)$ is the expectation of $f_{i}(\tau_{i},m_{i}({\tau}))$ conditionally on $(\tau_{l})_{l\neq i}$, we have
$\bigl{\langle}\bigl{(}f_{i}(\tau_{i},m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\bigl{(}f_{j}(\tau_{j},m_{j}^{(i)}({\tau}))-L_{j}^{(i)}({\tau})\bigr{)}\bigr{\rangle}=0,$
which implies
$\bigl{\langle}\bigl{(}f_{i}(\tau_{i},m_{i}({\tau}))-L_{i}({\tau})\bigr{)}\bigl{(}f_{j}(\tau_{j},m_{j}({\tau}))-L_{j}({\tau})\bigr{)}\bigr{\rangle}$
$=\bigl{\langle}X_{i}(\tau)\big{(}f_{j}(\tau_{j},m_{j}({\tau}))-f_{j}(\tau_{j},m_{j}^{(i)}({\tau}))\big{)}\bigr{\rangle}$ (14)
$+\bigl{\langle}X_{i}(\tau)c_{ij}({\tau})\bigr{\rangle},$ (15)